# Forecasting Communication Derailments Through Conversation Generation

### v1

@@ Abstract
Forecasting communication derailment can be useful in real-world settings such as online content moderation, conflict resolution, and business negotiations. However, despite language models’ success at identifying offensive speech present in conversations, they struggle to forecast future communication derailments. In contrast to prior work that predicts conversation outcomes solely based on the past conversation history, our approach samples multiple future conversation trajectories conditioned on existing conversation history using a fine-tuned LLM. It predicts the communication outcome based on the consensus of these trajectories. We also experimented with leveraging socio-linguistic attributes, which reflect turn-level conversation dynamics, as guidance when generating future conversations. Our method of future conversation trajectories surpasses state-of-the-art results on English communication derailment prediction benchmarks and demonstrates significant accuracy gains in ablation studies.
Forecasting Communication Derailments Through Conversation Generation
Yunfan Zhang1, Kathleen McKeown1, Smaranda Muresan1, 1Columbia University {yunfan, smara, kathy}@cs.columbia.edu
@@ 1 Introduction
Predicting future communication derailments from ongoing conversations has a wide range of real-world applications. For instance, in online moderation, the ability to forecast if a discussion thread might devolve into offensive exchanges allows moderators to intervene preemptively. Similarly, in conflict resolution, political hearings, and business negotiations, a system that warns participants of impending conflicts could help mitigate disputes before they escalate.
Despite its utility, predicting future communication derailments presents significant challenges. As shown in examples in Figure 1, unlike the detection of offensive conversational turns Warner and Hirschberg (2012); Poletto et al. (2021); Fortuna and Nunes (2018); Nobata et al. (2016), which focuses on identifying harmful speech within a given conversation transcript, our task requires forecasting whether offensive turns will occur later in the conversation. This makes our task significantly more challenging: the model has to predict potential future derailments based on an otherwise benign conversation history, and this demands a nuanced understanding of the conversation’s progression. Our task is further complicated by the inherently diverse nature of human interactions: given the same conversational history, there are multiple possible future trajectories, each with varying topics, styles, and tones.
To address these challenges, we propose a novel approach that forecasts future derailments by generating potential continuations of the given conversation. We define conversation derailments as conversations that end with offensive speech or ad hominem attacks. Our approach is based on the observation that while Large Language Models (LLMs) may not be effective classifiers for forecasting communication breakdowns directly, they excel at generating conversation continuations. Thus, we first fine-tune an LLM on modeling online conversations. We then sample multiple conversation continuations from this fine-tuned LLM, conditioned on the existing conversation history. We feed each generated continuation along with the existing conversation history to a binary conversation derailment classifier and obtain multiple predictions. We determine the final conversation outcome by taking a majority vote of these individual predictions. By randomly sampling multiple potential future conversations, we approximate the probability distribution of possible trajectories, thus helping us predict whether a derailment is likely.
@@ 2 Problem Statement and Motivation
Methods Acc. Prec. Rec. F-1 Acc. Prec. Rec. F-1
@@
Table 1: Accuracy, precision, recall, and F-1 scores on CGA-Wiki and BNC. Modern Language Models (LMs) can easily identify offensive speech present in the conversation. However, we notice a significant drop in accuracy when only the benign speech is given and the model is trained to forecast offensive speech in future exchanges.
@@
These findings emphasize a key challenge in this task: while LMs excel at identifying offensive speech present in the transcript, they struggle significantly when predicting upcoming derailments based solely on benign conversation history. This performance gap suggests that, in the absence of overt offensive speech, LMs have difficulty discerning the subtle conversational cues that may indicate a future shift toward harmful communication.
@@ 3 Methodology
Figure 2 shows an overview of our approach. We employ a fine-tuned LLM to generate multiple future continuations given the existing conversation history. We then train a dedicated classifier to assess the conversation outcome by analyzing both the existing and newly generated conversations. We then determine the final outcome based on the majority vote of these individual predictions. We also study whether social orientation labels (see Section 3.4) can be used to guide the text generation and conversation derailment classification.
@@ 3.1 Fine-tuning Conversation LLMs
During training, the first $k$ turns and the corresponding social orientation labels $\{(s_{1},c_{1}),\dots,(s_{k},c_{k})\}$ (if used) are used as the input context, denoted as $\mathbf{x}_{j}$. The model is trained to jointly predict the subsequent sequence of social orientation labels and conversation turns $\{(s_{k+1},c_{k+1}),\dots,(s_{n},c_{n})\}$, which forms $\mathbf{y}_{j}$. We experiment with both conditions: one where the social orientation labels are included and one where they are omitted. The training objective is:
@@ 3.3 Inference Time Majority Voting
At inference time, given a conversation history, we generate $L$ hypothetical conversation continuations using our fine-tuned conversation generation LLM $f_{\psi}$. Each generated continuation is appended to the existing conversation history and then classified by conversation outcome classifier $f_{\phi}$, resulting in $L$ individual outcome predictions. The final conversation outcome is determined by taking the majority vote across these $L$ predictions.
@@ 3.4 Social Orientation Labels
Building on the methodology outlined in Morrill et al. (2024), we annotate a set of social orientation labels on a turn-by-turn basis using GPT-4o. We extend over Morrill et al. (2024) by decoupling the axes and prompting GPT-4o to consider labels for each axis independently. Each turn is assigned one label from each of the four psychological dimensions. Formally, given an LLM $f_{\theta}$, a few-shot prompt $p$, and a conversation $C=\{c_{1},\dots,c_{n}\}$, the corresponding set of social orientation labels $S=\{s_{1},\dots,s_{n}\}$ is computed as follows:
@@ 4 Experiments
We evaluate our conversation derailment prediction method on two datasets: Conversation Gone Awry Wiki Split Zhang et al. (2018); Chang and Danescu-Niculescu-Mizil (2019) and the Before Naming Calling dataset Habernal et al. (2018). We describe the datasets and our experiment details below.
@@ 4.1 Datasets
is a communication derailment dataset derived from the discussion history between Wikipedia editors. We use both the original dataset Zhang et al. (2018) and the additional samples annotated in Chang and Danescu-Niculescu-Mizil (2019), resulting in a total of 4,188 samples. After excluding section headers without actual conversation turns, each sample comprises 3 to 19 turns, with a median of 6 turns. The first $n-1$ turns in each sample are benign, while the final turn is either benign or offensive with equal probability.
Following prior work Altarawneh et al. (2023); Kementchedjhieva and Søgaard (2021); Yuan and Singh (2023); Morrill et al. (2024), for the primary results in Section 5 and Table 2, we treat the first $k=n-1$ turns as the conversation history and input to our model. We then predict whether the last turn will be benign or offensive.
@@
Habernal et al. (2018) is a communication derailment dataset sourced from posts and replies in the Reddit r/changemyview community, containing 2,582 samples. Each sample includes $n=4$ consecutive conversation turns from the same comment-reply thread. The first $k=3$ turns are benign, whereas the 4th turn is either a post up-voted by the original poster ("successfully changed their view") or flagged for containing ad hominem attacks by Reddit moderators, with equal probability. We are aware that Zhang et al. (2018) also compiled a conversation derailment dataset (CGA-CMV) using comments from Reddit r/changemyview community; however, we prefer BNC over CGA-CMV as the BNC dataset contains less false-negatives upon our manual examination.
@@ 4.2 Experiment Setup
After fine-tuning of Mistral-7B for conversation generation, we augment our training set with hypothetical future turns generated by our fine-tuned model, as described in Section 3.2. We generate $l=2$ hypothetical continuations [^1] We choose $l=2$ and $L=5$ to balance computational cost with prediction accuracy. Increasing $l$ and $L$ enhances the ability to approximate future conversational trajectories and therefore improves prediction accuracy, at the cost of computational resources. We offer an ablation study on the effect of $L$ in Appendix A.2. for each training sample using the hyper-parameters listed in Table 6. We then fine-tune both a BART-Base model and a Mistral-7B-Base model to classify whether a derailment occurs in the transcript. We use the two synthetic conversation continuations along with the real future conversation turns as the input, and the gold derailment labels provided by the datasets as the target label in fine-tuning. The training hyper-parameters for the BART classifier are provided in Table 7 in Appendix A.4, and those for the Mistral classifier are in Table 8.
@@
We follow the inference protocol detailed in Section 3.3. For each test sample, we generate $L=5$ conversation continuations [^1] We choose $l=2$ and $L=5$ to balance computational cost with prediction accuracy. Increasing $l$ and $L$ enhances the ability to approximate future conversational trajectories and therefore improves prediction accuracy, at the cost of computational resources. We offer an ablation study on the effect of $L$ in Appendix A.2. using our LLM, with hyper-parameters specified in Table 6. Each of the 5 generated continuations is concatenated with the original conversation history and passed to our LM classifier, resulting in 5 individual predictions. The final prediction for communication derailment is based on the majority vote among these 5 predictions.
@@
We evaluate the performance of our approaches on accuracy, precision, recall, and F-1. We consider the presence offensive or ad hominem speech as the positive case for the purpose of calculating precision and recall.
@@ 5 Results and Analysis
Table 2: Accuracy, precision, recall, and F1 scores on CGA-Wiki and BNC. SFT stands for supervised fine-tuning, SO stands for social orientation labels, and G stands for generation. $\star$ denotes statistically significant (t-test $p<0.1$) accuracy gains compared to the baselines. $\bullet$ denotes statistically significant (t-test $p<0.1$) accuracy gains compared to the best-performing state of the art. On both datasets, methods that leverage conversation generation consistently outperforms the state-of-the-art methods and the baselines. Social orientation labels further improve performance on BNC when paired with conversation generation.
@@ 5.1 Comparisons with the State of the Art
As shown in Table 2, on the CGA-Wiki dataset, our best-performing model, the Mistral classifier with conversation generation (Mistral + G), surpasses the best previous fine-tuned models Altarawneh et al. (2023) in accuracy, precision, and F1 score, while slightly trailing in recall. It also outperforms GPT-4o few-shot in accuracy, precision, and F1 score. While our model falls short of GPT-4o in recall, GPT-4o’s high recall is artificially inflated due to its tendency to over-predict derailment. Despite balanced class labels (50%) in the datasets and few-shot examples, GPT-4o classifies 73.6% of conversations as derailments, leading to an inflated recall.
@@ 5.3 Impact of Social Orientation Labels
As in Table 2, we found mixed results on the effect of social orientation labels, depending on the dataset. The addition of social orientation labels enhanced the prediction of communication derailment on the BNC dataset. For both Mistral + G + SO and BART + G + SO, adding social orientation labels on top of conversation generation improved accuracy, recall, and F1 score, though there is a slight decrease in precision. However, incorporating social orientation labels did not enhance conversation derailment detection accuracy on CGA-Wiki dataset.
@@ 5.4 Impact of Generating Multiple Future Turns
Methods Category Acc. Prec. Rec. F-1 Acc. Prec. Rec. F-1
@@
Table 3: Accuracy, precision, recall, and F-1 scores on CGA-Wiki, but with only the first 2 or 4 turns as inputs to the model. SFT stands for supervised fine-tuning, and G stands for conversation generation. BART classifier with future conversation generation (BART + G) consistently achieves the highest accuracy in both first 2 turns and first 4 turns settings.
We investigate the impact of generating multiple future turns on communication derailment prediction. Instead of stipulating the number of future turns to generate, which could be considered leaking future information, we restrict the input context to the first $k=2$ and first $k=4$ turns, and then allow the model to generate multiple subsequent turns until it determines the conversation is complete. We observe that with the first 2 turns, the model generates a median of 3 more turns, while with the first 4 turns, it generates a median of 2.
As shown in Table 3, performance declines in both the "first 2 turns" and "first 4 turns" scenarios compared to the full $n-1$ turn input. This result is expected, as the model operates with reduced conversational context and needs to predict further into the future. However, in both cases, the method employing the BART classifier with future conversation generation (BART + G) maintained the highest accuracy. While the improvement is more modest (approximately 1%) compared to the full $n-1$ turn setting, the prediction-through-generation approach still proves beneficial.
@@ 6.1 Offensive Speech Detection
6.2 Predicting Communication Derailment
Several studies have addressed the challenge of predicting communication derailment using existing conversation history. Zhang et al. (2018); Chang and Danescu-Niculescu-Mizil (2019) created the CGA dataset by manually identifying ad hominem comments within Wikipedia editor discussions. Habernal et al. (2018) compiled BNC dataset by using the moderation results and upvotes from comments in Reddit r/changemyview community. Kementchedjhieva and Søgaard (2021) proposed adopting dynamic forecast window and pretrained LMs to predict derailments on CGA. Yuan and Singh (2023) applies a hierarchical transformer-based framework on CGA. Morrill et al. (2024) improved performance on the CGA by utilizing turn-level social orientation labels annotated by GPT-4. Hua et al. (2024) suggested employing conversation-level natural language summaries generated by GPT-4 to improve communication derailment prediction performance. Altarawneh et al. (2023) proposed using graph neural networks for predicting derailments on CGA.

### v2

# Forecasting Conversation Derailments Through Generation

@@ Abstract
Forecasting conversation derailment can be useful in real-world settings such as online content moderation, conflict resolution, and business negotiations. However, despite language models’ success at identifying offensive speech present in conversations, they struggle to forecast future conversation derailments. In contrast to prior work that predicts conversation outcomes solely based on the past conversation history, our approach samples multiple future conversation trajectories conditioned on existing conversation history using a fine-tuned LLM. It predicts the conversation outcome based on the consensus of these trajectories. We also experimented with leveraging socio-linguistic attributes, which reflect turn-level conversation dynamics, as guidance when generating future conversations. Our method of future conversation trajectories surpasses state-of-the-art results on English conversation derailment prediction benchmarks and demonstrates significant accuracy gains in ablation studies.
Forecasting Conversation Derailments Through Generation
Yunfan Zhang1, Kathleen McKeown1, Smaranda Muresan1,2 1Columbia University  2Barnard College yunfan.z@columbia.edukathy@cs.columbia.edusmara@columbia.edu
@@ 1 Introduction
Predicting future derailments from ongoing conversations has a wide range of real-world applications. For instance, in online moderation, the ability to forecast if a discussion thread might devolve into offensive exchanges allows moderators to intervene preemptively. Similarly, in conflict resolution, political hearings, and business negotiations, a system that warns participants of impending conflicts could help mitigate disputes before they escalate.
Despite its utility, predicting future conversation derailments presents significant challenges. As shown in examples in Figure 1, unlike the detection of offensive conversational turns Warner and Hirschberg (2012); Poletto et al. (2021); Fortuna and Nunes (2018); Nobata et al. (2016), which focuses on identifying harmful speech within a given conversation transcript, our task requires forecasting whether offensive turns will occur later in the conversation. This makes our task significantly more challenging: the model has to predict potential future derailments based on an otherwise benign conversation history, and this demands a nuanced understanding of the conversation’s progression. Our task is further complicated by the inherently diverse nature of human interactions: given the same conversational history, there are multiple possible future trajectories, each with varying topics, styles, and tones.
To address these challenges, we propose a novel approach that forecasts future derailments by generating potential continuations of the given conversation. We define conversation derailments as conversations that end with offensive speech or ad hominem attacks, with dataset-specific criteria and annotation procedures explained in Section 4.1. Our approach is based on the observation that while Large Language Models (LLMs) may not be effective classifiers for forecasting conversation breakdowns directly, they excel at generating conversation continuations. Thus, we adopt a generate-then-predict approach for forecasting conversation derailments. We first fine-tune an LLM on modeling online conversations. We then sample multiple conversation continuations from this fine-tuned LLM, conditioned on the existing conversation history. We feed each generated continuation along with the existing conversation history to a binary conversation derailment classifier and obtain multiple predictions. We determine the final conversation outcome by taking a majority vote of these individual predictions. By sampling multiple plausible continuations and aggregating the individual predictions with majority vote, we reduce the variability of stochastic LLM outputs, leading to more robust and accurate forecasting of conversation derailments.
@@ 2 Problem Statement and Motivation
Methods Acc. Prec. Rec. F1 Acc. Prec. Rec. F1
@@
Table 1: Accuracy, precision, recall, and F1 scores on CGA-Wiki and BNC. Modern Language Models (LMs) can easily identify offensive speech present in the conversation. However, we notice a significant drop in accuracy when only the benign speech is given and the model is trained to forecast offensive speech in future exchanges.
@@
These findings emphasize a key challenge in this task: while LMs excel at identifying offensive speech present in the transcript, they struggle significantly when predicting upcoming derailments based solely on benign conversation history. This performance gap suggests that, in the absence of overt offensive speech, LMs have difficulty discerning the subtle conversational cues that may indicate a future shift toward derailments.
@@ 3 Methodology
Figure 2 shows an overview of our approach. We employ a fine-tuned LLM to generate multiple future continuations given the existing conversation history. We then train a dedicated classifier to assess the conversation outcome by analyzing both the existing and newly generated conversations. We determine the final outcome based on the majority vote of these individual predictions. We also study whether social orientation labels (see Section 3.4) can be used to guide the text generation and conversation derailment classification.
@@ 3.1 Fine-tuning Conversation LLMs
During training, the first $k$ turns and the optional social orientation labels $\{(s_{1},c_{1}),\dots,(s_{k},c_{k})\}$ are used as the input context, denoted as $\mathbf{x}_{j}$. The model is trained to jointly predict the subsequent sequence of social orientation labels (if used) and conversation turns $\{(s_{k+1},c_{k+1}),\dots,(s_{n},c_{n})\}$, which forms $\mathbf{y}_{j}$. The training objective is:
@@ 3.3 Inference Time Majority Voting
At inference time, given a conversation history, we generate $L$ hypothetical conversation continuations using our fine-tuned conversation generation LLM $f_{\psi}$. Each generated continuation is appended to the existing conversation history and then classified by conversation outcome classifier $f_{\phi}$, resulting in $L$ individual outcome predictions. The final conversation outcome is determined by taking the majority vote across these $L$ predictions. This procedure compensates for the inherent randomness in LLM decoding, thereby reducing the influence of outlier generations and improving the robustness and accuracy of the final prediction.
@@ 3.4 Social Orientation Labels
Building on the methodology outlined in Morrill et al. (2024), we annotate a set of social orientation labels on a turn-by-turn basis using GPT-4o. We extend Morrill et al. (2024) by decoupling the axes and prompting GPT-4o to consider labels for each axis independently. Each turn is assigned one label from each of the four psychological dimensions. Formally, given an LLM $f_{\theta}$, a few-shot prompt $p$, and a conversation $C=\{c_{1},\dots,c_{n}\}$, the corresponding set of social orientation labels $S=\{s_{1},\dots,s_{n}\}$ is computed as follows:
@@ 4 Experiments
We evaluate our conversation derailment prediction method on two datasets: Conversation Gone Awry Wiki Split Zhang et al. (2018); Chang and Danescu-Niculescu-Mizil (2019) and the Before Name Calling dataset Habernal et al. (2018). We describe the datasets and our experiment details below.
@@ 4.1 Datasets
is a conversation derailment dataset derived from the discussion history between Wikipedia editors. In this dataset, derailments are defined as personal attacks or ad hominem speech. These labels were provided by paid human annotators. We use both the original dataset Zhang et al. (2018) and the additional samples annotated in Chang and Danescu-Niculescu-Mizil (2019), resulting in a total of 4,188 samples. After excluding section headers without actual conversation turns, each sample comprises 3 to 19 turns, with a median of 6 turns. The first n1n-1 turns in each sample are benign, while the final turn is either benign or offensive with equal probability.
Following prior work Altarawneh et al. (2023); Kementchedjhieva and Søgaard (2021); Yuan and Singh (2023); Morrill et al. (2024), for the primary results in Section 5 and Table 2, we treat the first k=n1k=n-1 turns as the conversation history and input to our model. We then predict whether the last turn will be benign or offensive.
@@
Habernal et al. (2018) is a conversation derailment dataset sourced from posts and replies in the Reddit r/changemyview community, containing 2,582 samples. Each sample includes $n=4$ consecutive conversation turns from the same comment-reply thread. The first $k=3$ turns are benign, while the 4th turn represents either a constructive or derailing outcome: a constructive outcome is one where the reply successfully changes the original poster’s view, as evidenced by an upvote from the OP; a derailing outcome is one where the reply is flagged by moderators for containing ad hominem or otherwise offensive content. These two outcomes are represented with equal probability in the dataset. We are aware that Zhang et al. (2018) also compiled a conversation derailment dataset (CGA-CMV) using comments from Reddit r/changemyview community; however, we prefer BNC over CGA-CMV as the BNC dataset contains fewer false-negatives upon our manual examination.
@@ 4.2 Experiment Setup
After fine-tuning Mistral-7B for conversation generation, we augment our training set with hypothetical future turns generated by our fine-tuned model, as described in Section 3.2. We generate $l=2$ hypothetical continuations for each training sample, and then fine-tune both a BART-Base model and a Mistral-7B-Base model to classify whether a derailment occurs in the transcript. We use the two synthetic conversation continuations along with the real future conversation turns as the input, and the gold derailment labels provided by the datasets as the target label in fine-tuning. The training hyper-parameters for the BART classifier are provided in Table 7 in Appendix A.4, and those for the Mistral classifier are in Table 8.
@@
We follow the inference strategy outlined in Section 3.3. To determine $L$, the number of continuations per sample, we conduct an ablation study on the CGA-Wiki and BNC validation sets, as detailed in Appendix A.2 and Table 4. We vary $L$ from 1 to 15 and observe that increasing $L$ from 1 to 5 yields a 1-2% improvement in accuracy. Beyond $L=5$, gains are marginal (0.1-0.4%). Based on this, we fix $L=5$ for all experiments to strike a balance between prediction accuracy and computational overhead.
@@
We evaluate the performance of our approaches on accuracy, precision, recall, and F-1. We consider the presence of offensive or ad hominem speech as the positive case for the purpose of calculating precision and recall.
@@ 5 Results and Analysis
Table 2: Accuracy, precision, recall, and F1 scores on CGA-Wiki and BNC. SFT stands for supervised fine-tuning, SO stands for social orientation labels, and G stands for generation. $\star$ denotes statistically significant (z-test $p<0.1$) accuracy gains compared to the baselines. $\bullet$ denotes statistically significant (z-test $p<0.1$) accuracy gains compared to the best-performing state of the art. On both datasets, methods that leverage conversation generation consistently outperforms the state-of-the-art methods and the baselines. Social orientation labels further improve performance on BNC when paired with conversation generation.
@@ 5.1 Comparisons with the State of the Art
As shown in Table 2, on the CGA-Wiki dataset, our best-performing model, the Mistral classifier with conversation generation (Mistral + G), surpasses the best previous fine-tuned models in accuracy, precision, and F1 score, while slightly trailing Kementchedjhieva and Søgaard (2021) in recall. It also outperforms GPT-4o few-shot in accuracy, precision, and F1 score. While our model falls short of GPT-4o in recall, GPT-4o’s high recall is artificially inflated due to its tendency to over-predict derailment. Despite balanced class labels (50%) in the datasets and few-shot examples, GPT-4o classifies 73.6% of conversations as derailments, leading to an inflated recall.
@@ 5.3 Impact of Social Orientation Labels
As in Table 2, we found mixed results on the effect of social orientation labels, depending on the dataset. The addition of social orientation labels enhanced the prediction of conversation derailment on the BNC dataset. For both Mistral + G + SO and BART + G + SO, adding social orientation labels on top of conversation generation improved accuracy, recall, and F1 score, though there is a slight decrease in precision. However, incorporating social orientation labels did not enhance conversation derailment detection accuracy on CGA-Wiki dataset.
@@ 5.4 Impact of Generating Multiple Future Turns
Methods Category Acc. Prec. Rec. F1 Acc. Prec. Rec. F1
@@
Table 3: Accuracy, precision, recall, and F1 scores on CGA-Wiki, but with only the first 2 or 4 turns as inputs to the model. SFT stands for supervised fine-tuning, and G stands for conversation generation. BART classifier with future conversation generation (BART + G) consistently achieves the highest accuracy in both first 2 turns and first 4 turns settings.
We investigate the impact of generating multiple future turns on conversation derailment prediction. Instead of stipulating the number of future turns to generate, which could be considered leaking future information, we restrict the input context to the first $k=2$ and first $k=4$ turns, and then allow the model to generate multiple subsequent turns until it determines the conversation is complete. We observe that with the first 2 turns, the model generates a median of 3 more turns, while with the first 4 turns, it generates a median of 2.
As shown in Table 3, performance declines in both the "first 2 turns" and "first 4 turns" scenarios compared to the full n1n-1 turn input. This result is expected, as the model operates with reduced conversational context and needs to predict further into the future. However, in both cases, the method employing the BART classifier with future conversation generation (BART + G) maintained the highest accuracy. While the improvement is more modest (approximately 1%) compared to the full n1n-1 turn setting, the prediction-through-generation approach still proves beneficial.
5.5 Diversity of Multiple Conversation Continuations
We calculate the BLEU score between different continuations given the same conversation history to evaluate the diversity of generated conversation continuations. More specifically, for each set of $L=5$ continuations generated for the same conversation, we compute BLEU score by treating one continuation as the hypothesis and the remaining four as references. We repeat this for all possible (4+1) combinations and report the average BLEU score. Lower scores indicate higher diversity among the generated continuations. We focus on the top-performing methods on each dataset: Mistral + G on CGA-Wiki and Mistral + SO + G on BNC.
We found our technique yields highly diverse conversation continuations, with average BLEU scores between inputs at 0.034 for Mistral + G on CGA-Wiki and 0.046 for Mistral + SO + G on BNC. These low scores indicate that the generated continuations are substantially dissimilar from each other, allowing our method to forecast a broad and varied set of plausible conversational trajectories.
@@ 6.1 Offensive Speech Detection
6.2 Predicting Conversation Derailment
Several studies address the challenge of predicting conversation derailment using existing conversation history. Zhang et al. (2018); Chang and Danescu-Niculescu-Mizil (2019) create the CGA dataset by manually identifying ad hominem comments within Wikipedia editor discussions. Habernal et al. (2018) compile BNC dataset by using moderation results and upvotes from comments in Reddit r/changemyview community. Kementchedjhieva and Søgaard (2021) propose adopting dynamic forecast window and pretrained LMs to predict derailments on CGA. Yuan and Singh (2023) apply a hierarchical transformer-based framework on CGA. Morrill et al. (2024) improve performance on the CGA by utilizing turn-level social orientation labels annotated by GPT-4. Hua et al. (2024) suggest employing conversation-level natural language summaries generated by GPT-4 to improve conversation derailment prediction performance. Altarawneh et al. (2023) propose using graph neural networks for predicting derailments on CGA. Yuzbashyan et al. (2025) explore using LLMs such as GPT-4 and Llama 3 to create synthetic training data for CGA.
In contrast to these methods, which rely on existing conversation history only, our approach generates plausible future turns and bases its prediction on both the existing dialogue and these potential continuations. This allows our derailment classifier to consider how the conversation might unfold in its decision.